A body floats with 2/3 of its volume outside water and 3/4 of its volume outside the other liquid . The density of the liquid is

Grade:11

1 Answers

Arun
6 years ago
Dear Ashish
 
Volume = V
density= d
density of water =q = 1000 kg/m3
weight in air = w= Vgd
Buoyant force= (V/3)qg

density of other liquid =p
volume of body inside water= V- 3V/4= V/4
buoyant force = (V/4)pg

as in both situation the body will float so, 
(V/3)qg = (V/4)pg
p = 4q/ 3
p = 4 x 1000/3
p= 4 /3 x 10³
 
Regards
Arun (askIITians forum expert)